OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. Find w if P = 68 and l = 10. A. W = 25 B. W = 15 C. W = 28 D. W = 24

OpenStudy (anonymous):

Answer can be found by substitution. P=68 and l=10 Put these into P=2l+2w 68=2(10)+2w 68=20+2w 2w+20=68 2w+20-20=68-20 2w=48 2w/2=48/2 w=48/2 =24

OpenStudy (lgbasallote):

detailed explanation @CoCoTsoi good job ;D (Y) <---that's a thumbs up

OpenStudy (anonymous):

Thanks, lgba! I will keep going :D